OpenStudy (anonymous):

Which of the following is false? A. 5 + 7 = 7 + 5 B. 2003-01-04-00-00_files/i0050000.jpg C. 2003-01-04-00-00_files/i0050001.jpg D. 2003-01-04-00-00_files/i0050002.jpg

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

well it is not a and b,c,and d are not showing correctly

OpenStudy (anonymous):

Well, the first is true because 5 +7 = 7 +5. That is a common Property of Addition. I cannot see b, c, and d, because it lists a .jpg file name, not the actual pictures.

OpenStudy (texaschic101):

D is not correct

OpenStudy (anonymous):

so whats the right answer

OpenStudy (texaschic101):

D is false

OpenStudy (texaschic101):

this is what D says....6x^3 - 9x = 9x - 6x^3...this is false

OpenStudy (anonymous):

D is false.

OpenStudy (anonymous):

False

OpenStudy (anonymous):

thank yall so much